# Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions

### v1

@@ 4.1 Does SD prompting improve alignment?
To study RQ1, we compare SD vs. base prompts. What is the effect of adding SD information on opinion alignment? Note that while prior work studies this question only for majority-voted responses Beck et al. (2024); Hu and Collier (2024), we instead look at the effect of adding SD information when evaluating distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets. The results for all models are shown in Appendix Table 5.
For both base and SD prompts, LLM-generated distributions are generally best with verbalized elicitation (Table 1), though with exceptions (e.g., OLMo-Instruct on the OQA dataset). As shown in the table, prompting with SD information does not necessarily increase opinion alignment with human responses. In fact, it is often comparable or even lower than prompting without any SD information. The degree of effectiveness in creating distributions that align with human opinion appears to be model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024).
@@ 4.2 Can we calibrate LLM distributions?
We have seen that prompting LLMs with SD information does not consistently increase opinion alignment with human responses. Since LLM-generated distributions have fairly low opinion alignment, with RQ2, we question whether we can calibrate these distributions to better align them with human response distributions on these survey datasets.
@@
The results after regression, compared to pre-regression results, are shown in Table 1, with results for all models in Appendix Table 5. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, and by an average of 16.3%. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment decreases by an average of 4.1%. These results suggest that LLM-generated distributions for predicting human responses are somewhat uncalibrated, and a simple supervised regression can lead to higher opinion alignment.
@@
After calibration, opinion alignment in each setting also has much lower variance. Calibrated distributions tend to have a much lower standard deviation; standard deviation is lower in 87.2% of dataset-LLM-elicitation settings, and is on average 1.62 times lower. That lower standard deviation leads to higher opinion alignment provides more evidence that LLMs might be exaggerating differences in opinion distributions between demographic groups Cheng et al. (2023a, b), and that calibration could be helpful in mitigating this.
@@
In addition, we find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. We see that the standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. While prior work Beck et al. (2024); Hu and Collier (2024) and our own previous results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
@@ 5 How does alignment vary across SDs?
So far, we have studied how well LLMs can predict responses from SD groups aggregated across all groups, with the goal of comparing different models and distribution elicitation methods. However, a particular model or method might be more aligned with some sociodemographics over others. With RQ3, we seek to understand the differences at the more granular demographic level.
@@
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it performed the best with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. We choose 5 out of 15 LLMs; the most recent/powerful from each model family: Claude 3.5 v2, Llama 3.2 90B, Mistral large, OLMo-2 7B Instruct, and Qwen 2.5 72B. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 2. See Appendix A.4 for alignment for other demographics.
Findings (RQ3):
(1) Calibration produces distributions more aligned with some demographics over others; (2) Claude models are more highly aligned with OQA dataset SD groups; (3) different models are better aligned with specific demographics.
5.1 How does calibration affect alignment?
@@
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all sociodemographic groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more aligned for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. As expected, calibration increases alignment for some demographics and reduces alignment for others. We find that it tends to increase alignment of SD groups that were less aligned with base-prompted distributions, such as Central Africa, Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment is decreased for very high aligned demographic groups, such as Aus/NZ and Northern Europe with Llama 90B.
5.2 How does alignment vary across LLMs?
@@
Next, we compare the opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Interestingly, Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less of an important factor in determining the answer to the survey questions for those with middle income.
@@ 6 Does supervision work with less data?
Although we used 80% of our data as supervision for our regression models to calibrate the LLM-generated distributions, we study how opinion alignment varies with smaller amounts of supervised data to explore RQ4.
Findings (RQ4):
(1) As few as 5 training examples to the regression model can suffice to achieve close to minimal MSE; (2) on average, degradation in alignment for individual demographics is close to zero at 5 examples; (3) demographics with the largest opinion alignment degradation is dependent on base vs. SD prompts; and (4) regression models generalize fairly well to unseen datasets.
6.1 How much supervised data do you need?
@@
We evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset. We average over 10 different random samples for each training data size. We find that as few as 5 examples (with 4 answer choices per example) can be enough to achieve close to the minimal MSE for any particular dataset-LLM-elicitation setting. We plot Mean Squared Error (MSE) over training data size in Appendix Figure 4 showing that MSE usually converges at 5 examples, though this is model and dataset dependent.
6.2 How are individual SD groups affected?
@@
How does using a small set of random examples affect alignment of individual demographics? Although degradation on average is close to zero, individual demographics are affected differently. Table 3 shows those most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Looking at absolute values, using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).

### v2

@@ 4.1 Does SD prompting improve alignment?
For RQ1, we compare SD vs. base prompts to study the effect of adding SD information on opinion alignment. Note that while prior work studies this question only for majority-voted responses Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024), we investigate the effect of adding SD information when comparing distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets.
For base and SD prompts, LLM distributions are generally most aligned with verbalized elicitation (Table 1), with exceptions (e.g., on the OQA dataset). As shown in the table, prompting with SD information often leads to comparable or even lower opinion alignment than prompting without any SD information, although this is model, dataset, and elicitation method dependent.
@@ 4.2 Can we calibrate LLM distributions?
We have seen that prompting LLMs with SD information does not consistently improve opinion alignment with human responses. We next investigate for RQ2 whether calibrating these distributions can more accurately and consistently align them with human response distributions.
@@
The results after regression, compared to pre-regression results, are shown in Table 1. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, improving by an average of 16.3%, where we average the delta between uncalibrated and calibrated opinion alignment across dataset-LLM-elicitation method settings. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment in these settings decreases by an average of 4.1%. These results suggest that LLM-generated distributions are somewhat uncalibrated, and a simple supervised regression can improve alignment.
| Model | Base prompt | Sociodemographic prompt | |||||||||||||||||
| WGM | Claude | 66.2 | 65.6 | 85.2 | 59.5 | 60.7 | 85.1 | 89.3 | 88.8 | 87.7 | 65.4 | 61.7 | 84.4 | 61.4 | 58.6 | 84.2 | 89.0 | 79.7 | 89.3 |
| OLMo | 59.6 | 62.9 | 81.5 | 67.7 | 62.7 | 80.4 | 59.8 | 86.1 | 85.1 | 62.3 | 64.0 | 82.5 | 64.5 | 59.8 | 82.7 | 69.8 | 81.8 | 84.6 | |
| OQA | Claude | 70.5 | 76.4 | 88.8 | 72.5 | 74.3 | 90.4 | 91.7 | 89.9 | 91.7 | 76.1 | 68.0 | 89.9 | 73.3 | 66.1 | 89.6 | 91.9 | 80.3 | 91.6 |
| OLMo | 72.6 | 79.3 | 89.0 | 72.3 | 77.0 | 88.4 | 65.4 | 85.2 | 79.9 | 72.8 | 77.1 | 88.4 | 70.5 | 76.3 | 88.6 | 68.5 | 82.9 | 81.5 | |
| WVS | Claude | 46.5 | 55.5 | 80.3 | 51.0 | 52.8 | 80.3 | 75.2 | 80.3 | 80.3 | 61.0 | 61.4 | 80.4 | 56.8 | 59.8 | 80.4 | 75.6 | 79.7 | 81.7 |
| OLMo | 75.3 | 76.9 | 80.3 | 74.9 | 73.5 | 80.3 | 86.6 | 86.5 | 86.5 | 58.3 | 72.7 | 79.2 | 60.0 | 69.4 | 77.2 | 86.0 | 85.7 | 89.8 | |
@@
After calibration, opinion alignment in each setting also has much lower variance. Calibrated distributions have a lower standard deviation in 87.2% of dataset-LLM-elicitation settings, and are 1.62 times lower on average. That lower standard deviation leads to higher opinion alignment provides further evidence that LLMs may exaggerate differences in opinion distributions between demographic groups Cheng et al. (2023a, b), and that calibration could help mitigate this.
@@
We also find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. Standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. Although there are some settings where uncalibrated distributions have a lower standard deviation than their calibrated counterparts, we emphasize that the variance between settings is much lower after calibration. While prior work Beck et al. (2024); Hu and Collier (2024) and our own results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
We do not find distinct patterns with model size within model family (e.g., Mistral small vs. Mistral large). Indeed, smaller models are sometimes more aligned than larger models, though across prompting methods, elicitation methods, and datasets, larger models are more aligned more consistently (see Appendix Table 7). However, we do find that chain of thought prompting (Section 4.3) appears to increase alignment for smaller models, but not for larger models. The smaller Llama models in particular often refused to answer questions (i.e., questions would hit safety restrictions for attempting to role-play as a particular identity group), or would not follow instructions in formatting output distributions. We were thus unable to extract distributions for all questions, leading to higher variance in alignment. See Appendix Table 17 for percent distributions extracted in each setting.
@@
4.3 The effect of chain of thought prompting
We compare CoT prompting with standard prompting + calibration for two models, the model that is most aligned on average (Claude 3.5 v2) and a small model (OLMo-2-7B Instruct). Our results are shown in Table 2. We find that for Claude, CoT does not consistently improve alignment over standard prompting, though it does improve in some settings. For base prompting with verbalized elicitation, alignment with CoT is not significantly different than with calibration for 2/3 datasets. For SD prompting, calibration creates significantly more aligned distributions for all datasets. For OLMo, we find that CoT improves alignment much more consistently, often leading to distributions that are just as aligned as, or better aligned than, calibrated distributions for the verbalized setting. Calibration is consistently more aligned than CoT for distributions elicited via paraphrase or self-random.
Upon qualitative inspection, we found that SD + CoT prompting Claude led to an “overthinking” effect Pu et al. (2025), where the model sometimes generated a distribution that exaggerated a particular difference with the ground truth. For example, the knowledge that “Men generally report higher confidence in scientific knowledge than women due to societal factors” to generate a distribution for the question “How much do you know about science?” led to the model producing a distribution that was more skewed than the ground truth. We hypothesize that this led to lower alignment for Claude (and would be similar for other larger, more capable reasoning models), but not for smaller models like OLMo 7B. Future work might study this further.
@@ 5 How does alignment vary across SDs?
So far, we have studied alignment with SD groups aggregated across groups, in order to compare different models and distribution elicitation methods. However, a particular model or method might be more aligned with some sociodemographics over others. With RQ3, we seek to understand the differences at the more granular demographic level.
@@
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it leads to the most aligned distributions with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 3. We show results for the same 5 LLMs used in Section 4. See Appendix A.3 for alignment for other demographics.
5.1 The effect of calibration
@@
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all SD groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more absolutely aligned than their uncalibrated counterparts for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. This includes all demographic categories, including those in Table 3 and Appendix A.3. The delta improvement across demographics is most significant for OLMo on the WGM and OQA datasets. As expected, calibration improves alignment for some demographics and reduces alignment for others. We find that it tends to improve alignment of SD groups that were less aligned with base-prompted distributions, such as Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment decreases for very high aligned demographic groups, such as Aus/NZ and Northern Europe.
| Model | Base prompt | Sociodemographic prompt | |||||||||||||||||
| WGM | Claude | 66.2 | 63.5 | 86.4 | 59.5 | 61.7 | 83.7 | 89.3 | 89.0 | 89.2 | 65.4 | 59.8 | 84.1 | 61.4 | 58.1 | 84.2 | 89.0 | 78.9 | 89.0 |
| Llama | 68.1 | 72.5 | 82.8 | 73.0 | 66.5 | 86.4 | 84.8 | 86.0 | 89.2 | 70.6 | 66.7 | 86.0 | 67.6 | 64.9 | 85.6 | 85.0 | 78.7 | 89.6 | |
| Mistral | 62.4 | 59.5 | 84.5 | 72.0 | 55.4 | 83.7 | 89.4 | 87.2 | 89.6 | 68.4 | 58.4 | 85.8 | 63.0 | 55.3 | 84.9 | 87.3 | 78.1 | 88.5 | |
| OLMo | 59.6 | 70.5 | 81.1 | 67.7 | 70.9 | 81.0 | 59.8 | 81.1 | 83.7 | 62.3 | 64.2 | 82.2 | 64.5 | 66.0 | 81.6 | 69.8 | 79.3 | 80.6 | |
| Qwen | 57.7 | 60.6 | 82.8 | 53.3 | 59.0 | 82.1 | 88.2 | 88.3 | 87.3 | 66.0 | 58.5 | 85.2 | 63.4 | 55.9 | 85.4 | 89.1 | 80.4 | 89.4 | |
| Average | 62.8 | 65.3 | 83.5 | 65.1 | 62.7 | 83.4 | 82.3 | 86.3 | 87.8 | 66.5 | 61.5 | 84.7 | 64.0 | 60.0 | 84.3 | 84.0 | 79.1 | 87.4 | |
| OQA | Claude | 70.5 | 77.9 | 88.4 | 72.5 | 75.7 | 89.6 | 91.7 | 91.5 | 91.5 | 76.1 | 68.1 | 89.7 | 73.3 | 66.9 | 89.7 | 91.9 | 79.7 | 92.3 |
| Llama | 79.3 | 77.6 | 89.7 | 75.3 | 76.6 | 89.6 | 86.8 | 88.8 | 88.6 | 79.2 | 75.0 | 90.0 | 76.1 | 75.5 | 89.8 | 83.4 | 80.5 | 87.1 | |
| Mistral | 79.5 | 77.4 | 89.6 | 75.3 | 74.0 | 89.0 | 85.0 | 89.7 | 86.9 | 75.8 | 70.5 | 89.8 | 72.4 | 69.9 | 87.5 | 83.8 | 80.7 | 87.8 | |
| OLMo | 72.6 | 77.0 | 88.9 | 72.3 | 74.6 | 85.9 | 65.4 | 81.2 | 77.4 | 72.8 | 72.3 | 88.6 | 70.5 | 74.4 | 88.4 | 68.5 | 78.8 | 86.6 | |
| Qwen | 73.9 | 75.9 | 90.1 | 67.0 | 74.4 | 89.3 | 88.4 | 88.9 | 87.2 | 74.4 | 71.4 | 90.2 | 71.3 | 69.3 | 89.3 | 89.2 | 81.0 | 87.8 | |
| Average | 75.2 | 77.2 | 89.3 | 72.5 | 75.1 | 88.7 | 83.5 | 88.0 | 86.3 | 75.7 | 71.5 | 89.7 | 72.7 | 71.2 | 88.9 | 83.4 | 80.1 | 88.3 | |
| WVS | Claude | 46.5 | 62.4 | 81.7 | 51.0 | 62.1 | 80.8 | 75.2 | 82.1 | 81.5 | 61.0 | 64.7 | 80.4 | 56.8 | 64.0 | 80.2 | 75.6 | 81.9 | 80.4 |
| Llama | 61.6 | 70.1 | 79.4 | 59.1 | 68.9 | 80.9 | 64.6 | 79.7 | 83.5 | 62.1 | 70.3 | 80.0 | 59.5 | 72.2 | 80.9 | 67.7 | 78.8 | 86.2 | |
| Mistral | 48.8 | 63.9 | 81.4 | 44.0 | 61.0 | 80.6 | 72.8 | 81.0 | 81.1 | 54.3 | 65.0 | 82.4 | 51.5 | 62.7 | 81.4 | 76.6 | 78.7 | 82.5 | |
| OLMo | 75.3 | 72.9 | 80.0 | 74.9 | 74.2 | 80.2 | 86.6 | 84.2 | 87.8 | 58.3 | 68.8 | 78.8 | 60.0 | 72.3 | 77.5 | 86.0 | 78.0 | 89.7 | |
| Qwen | 39.6 | 58.3 | 82.1 | 42.4 | 56.8 | 81.3 | 74.0 | 78.4 | 81.0 | 49.1 | 59.0 | 81.5 | 49.4 | 56.8 | 81.2 | 73.4 | 78.3 | 76.6 | |
| Average | 54.4 | 65.5 | 80.9 | 54.3 | 64.6 | 80.8 | 74.6 | 81.1 | 83.0 | 57.0 | 65.6 | 80.6 | 55.4 | 65.6 | 80.2 | 75.9 | 79.1 | 83.1 | |
5.2 Alignment variation across LLMs
@@
We now compare opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7.
In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less important in determining survey answers for those with middle income.
@@ 6 Does supervision work with less data?
Our results presented thus far used 80% of each dataset for training our supervised regression models to calibrate the LLM-generated distributions. We next investigate whether smaller amounts of supervision can suffice for opinion alignment (RQ4).
6.1 Amount of supervised data required
@@
As discussed in Section 3.6, we evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset, and average over 10 random samples for each size. We find that as few as 1-10 examples (with 4 answer choices per example) can suffice to closely match the minimal MSE for any particular dataset-LLM-elicitation setting. Figure 3, plotting Mean Squared Error (MSE) over training data size, shows that MSE usually converges between 1 and 10 examples, though this is model and dataset dependent.
6.2 Comparison with few-shot prompting
Given that the regression models appear to converge with so few examples, one might ask whether the regression is actually learning anything useful, or if we could achieve a similar improvement with few shot prompting instead. To investigate this, we 5-shot prompt the LLMs with the same examples given to the regression model in the 5 example calibration setting. We use a 5-shot prompt since this lies in the middle of how many examples were needed for convergence for the regression models. Whereas we averaged over 10 random samples in those experiments, here we use only 3 of the random samples because we found little variation. We average alignment over the 3 runs.
Table 4 compares few shot prompted () distributions to original and calibrated () distributions for base and SD prompts. While 5-shot prompted distributions are not as aligned as calibrated distributions in most settings, they are significantly more aligned in a few cases. However, calibration leads to distributions that are more aligned more consistently, across models, prompting methods, and distribution elicitation methods.
6.3 The effect on individual SD groups
@@
Although we see convergence with fewer examples, does using a small set of random examples affect alignment of individual demographics? We calibrate with a set of 5 examples (in the middle of the 1-10 range described above for MSE convergence). Although degradation on average is close to zero, individual demographics are affected differently. Table 5 shows the five SDs most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are historically less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + supervised finetuning (SFT), base + SFT + direct preference optimization (DPO), and the instruct model (base + SFT + DPO + reinforcement learning with verifiable rewards (RLVR)). The SFT step involves finetuning on instruction datasets and scaled synthetic data, the DPO step involves preference finetuning with DPO, and the RLVR step involves finetuning targeted at domains where prompts with verifiable rewards can be constructed, such as math OLMo et al. (2025). We use these four models out-of-the-box via Huggingface and study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it is typically most aligned.
Figure 4 plots opinion alignment for all OLMo models. Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets. OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to have less effect in this setting.
We suspect the increase in alignment on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS. The drop in alignment for uncalibrated distributions on WGM and OQA might indicate that OLMo’s post-training data is less aligned with some of the populations of these datasets. Overall, calibration leads to more consistent alignment, with and without SD prompting, regardless of post-training or dataset. This is in agreement with the rest of our results, where using calibration brings more consistency across settings.